Complete the table below. 1. Output VC ТC AVC AFC АТС МC 0 100 1 25 20 2 53.3 17.5 90 6 30 265 7 8 41.3 35.0 9 10 425 LO
Answer #1

TC = VC + FC

AVC = VC / Q

AFC = FC / Q

ATC = TC / Q

MC = Change in TC / Change in Q

Q VC TC AVC AFC ATC MC
0 100
1 25 125 25.0 100.0 125.0 25.0
2 45 145 22.5 50.0 72.5 20.0
3 59.9 159.9 20.0 33.3 53.3 14.9
4 70 170 17.5 25.0 42.5 10.1
5 90 190 18.0 20.0 38.0 20.0
6 120 220 20.0 16.7 36.7 30.0
7 165 265 23.6 14.3 37.9 45.0
8 230.4 330.4 28.8 12.5 41.3 65.4
9 315 415 35.0 11.1 46.1 84.6
10 425 525 42.5 10.0 52.5 110.0